# Recursive algorithms for computing Birkhoff interpolation polynomials

### v1

@@ Abstract
As a generalization of Hermite interpolation problem, Birkhoff interpolation is an important subject in numerical approximation. This paper generalizes the existing Generalized Recursive Polynomial Interpolation Algorithm (GRPIA) that is used to compute the Hermite interpolation polynomial. Based on the theory of the Schur complement and the Sylvester identity, the proposed recursive algorithms are applicable to a broader class of Birkhoff interpolation problems, where each interpolation condition is given by the composition of an evaluation functional and a differential polynomial. The approach incorporates a judgment condition to ensure the problem’s well-posedness and computes a lower-degree Newton-type interpolation basis (which is also a strongly proper interpolation basis) along with the corresponding interpolation polynomial. Compared with existing algorithms that rely on Gaussian elimination to compute the interpolation basis, our recursive approach significantly reduces the computational cost.
@@
pacs:
Mathematics Subject Classification (2010)]MSC 41A05 $\cdot$ MSC 65F $\cdot$ MSC 15A
@@ 1 Introduction
Schoenberg proposed the theory of the so-called incidence matrix and its regularity in [18], after which research on the Polya’s conditions has advanced rapidly. Polya’s conditions provide a necessary and sufficient criterion for the well-posedness of a two-point interpolation problem. In [20], Turán investigated the (0,2)-interpolation problem and proposed several open problems related to Birkhoff interpolation. Shi presented a consolidation of the established results concerning the regularity of both the (0,2)-interpolation problem and specific Birkhoff interpolation problems [19]. In [21], Xia et al. investigated the linearization of univariate Birkhoff rational interpolation and demonstrated the effectiveness of these rational methods through numerical experiments.
The Birkhoff multivariate interpolation scheme $(Z,S,E)$ consists of three components: the set of interpolation nodes $Z$, the interpolation space $S$, and the incidence matrix $E$. Atkinson proposed the Polya’s conditions for the incidence matrix corresponding to the multivariate Birkhoff interpolation problem in [1]. Bojanov and his colleagues analyzed the Birkhoff interpolation problem in the two-dimensional case, where the nodes are distributed in concentric circles [2]. In addition to establishing necessary and sufficient conditions for the regularity of an interpolation scheme $(S,E)$, Crainic derived sufficient conditions for its singularity and almost-everywhere regularity, as well as the equivalence among different types of such schemes and their auxiliary conditions [6]. In [7], Crainic investigated the two-dimensional Birkhoff interpolation problem where the interpolation conditions at rectangular lattice points are consistent, and the index sets corresponding to each condition are also distributed in a rectangular pattern. Furthermore, he derived the monomial basis for the interpolation problem in specific special cases. The algorithm proposed by Chai et al. in [4] allowed the computation of the minimal degree interpolation polynomial for multivariate interpolation problems. In [11], Kashpur considered the Hermite–Birkhoff interpolation problem for a nonlinear operator in Hilbert space.
@@
Recent work by Rhouni et al. has addressed the Birkhoff interpolation problem where the conditions involve only derivatives at the nodes. In [17], they proposed the RHBPIA algorithm to determine the interpolation polynomial. Specifically, by constructing a set of high-degree Newton-type basis that can be factorized, the interpolation polynomial for the Birkhoff interpolation problem can be derived. This paper considers a more general Birkhoff interpolation problem in which each interpolation condition is given by the composition of an evaluation functional and a differential polynomial. We propose two algorithms to compute a lower-degree interpolation basis and the corresponding interpolation polynomial.
The structure of this paper is as follows. Section 2 introduces fundamental concepts of the Birkhoff interpolation problem, including the incidence matrix, the Schur complement, and the Sylvester identity. Section 3 defines the concept of a strongly proper interpolation basis and derives a recursive interpolation formula for Birkhoff interpolation using the Schur complement. Section 4 proposes two algorithms for the univariate Birkhoff interpolation problem and provides numerical examples to demonstrate their performance. The final section provides conclusions.
@@ 2 Preliminary
We first introduce some relevant concepts related to Birkhoff interpolation problem. Throughout the paper, let $\mathbb{F}$ denote either the real field $\mathbb{R}$ or the complex field $\mathbb{C}$. $\mathbb{F}[x]$ stands for the polynomial ring over $\mathbb{F}$. Let $x_{0},x_{1},\dots,x_{n}$ be $n+1$ pairwise distinct interpolation nodes.
Definition 1.
[18] Let
$E=(e_{i,j}),i=0,1,\dots,n,j=0,1,\dots,N,$
be a matrix with $n+1$ rows and $N+1$ columns, where $e_{i,j}=0\text{ or }1$, and $\sum e_{i,j}=N$. This matrix is called an incidence matrix. No row of the matrix $E$ can consist entirely of zeros.
For any given incidence matrix $E$, we define the set of pairs
$\tilde{e}=\{(\beta,\alpha)\mid e_{\beta,\alpha}\in E~and~e_{\beta,\alpha}=1\}.$ (1)
The Birkhoff interpolation problem is to find a polynomial subspace $\mathcal{F}\subset\mathbb{F}[x]$ such that there exists a unique $p(x)\in\mathcal{F}$ satisfying the following interpolation conditions
@@
Here $\beta_{1},\beta_{2},\dots,\beta_{N}$ are the row indices of $E$ that correspond to the node numbers $0,1,\dots,n$, and $\alpha_{1},\alpha_{2},\dots,\alpha_{N}$ are the column indices of $E$ which correspond to the order of differentiation.
In the following text, for a given Birkhoff interpolation problem (or incidence matrix), we always sort the interpolation conditions in accordance with the above definition. Note that the interpolation conditions can be expressed by linear functionals, i.e., the composite operation of evaluation functionals and differential operators
@@
in which $T=[(\beta_{1},\alpha_{1}),(\beta_{2},\alpha_{2}),\dots,(\beta_{N},\alpha_{N})]$ is an NN-DOS.
@@
Next, we introduce the concept of Schur complement and some of its properties.
@@
where the submatrix $A$ is assumed to be square and non-singular. The Schur complement of $A$ in $M$, denoted by $(M/A)$, is defined as
$(M/A)=D-CA^{-1}B.$
This formula can also be expressed as
$(M/D)=A-BD^{-1}C.$ (3)
@@
[3] Let $D$ be a non-singular matrix and $M$ a square matrix. Then we have
$|M|=|D|\cdot|(M/D)|,$ (4)
where $|M|$ denotes the determinant of the matrix $M$.
@@
[5, 13]Assuming that $D$ is a non-singular matrix, we have
$\left(\left(\begin{array}[]{cc}A&B\\ C&D\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}D&C\\ B&A\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}B&A\\ D&C\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}C&D\\ A&B\\ \end{array}\right)/D\right).$ (5)
@@
[5, 13]Supposing that $E$ is a given linear operator, $D$ is a non-singular matrix, and both $E\ast A$ and $E\ast B$ are well-defined, then we have
$\left(\left(\begin{array}[]{cc}E*A&E*B\\ C&D\\ \end{array}\right)/D\right)=E*\left(\left(\begin{array}[]{cc}A&B\\ C&D\\ \end{array}\right)/D\right).$ (6)
@@
[13, 16](The Sylvester identity)Let $M$ be the matrix defined in Definition 4, and $K$ be a partitioned matrix as shown below
@@
$(K/M)=((K/A)/(M/A))=\left(\left(\begin{array}[]{cc}E&F\\ H&A\\ \end{array}\right)/A\right)-\left(\left(\begin{array}[]{cc}F&G\\ A&B\\ \end{array}\right)/A\right)(M/A)^{-1}\left(\left(\begin{array}[]{cc}H&A\\ L&C\\ \end{array}\right)/A\right).$ (7)
@@ 3 Main results
$p(x)=\sum_{i=0}^{N-1}a_{i}q_{i}(x),$ (8)
@@
where $Y_{N}=(y_{\beta_{1},\alpha_{1}},y_{\beta_{2},\alpha_{2}},\dots,y_{\beta_{N},\alpha_{N}})^{\textup{T}}$ and
@@
$p(x)=(q_{0}(x),q_{1}(x),\dots,q_{N-1}(x))\boldsymbol{a}=(q_{0}(x),q_{1}(x),\dots,q_{N-1}(x))V_{N-1}^{-1}Y_{N}.$ (11)
@@
where $\{q_{0}(x),q_{1}(x),\dots,q_{N-1}(x)\}$ is a strongly proper interpolation basis.
@@
$V_{k-1}:=\left(\begin{array}[]{ccccc}q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&&\vdots\\ q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right).$ (12)
@@
Let $\textup{span}\{q_{0}(x),q_{1}(x),\dots,q_{N-1}(x)\}$ be a strongly proper interpolation space for the Birkhoff interpolation problem (2). For any fixed $k=1,2,\dots,N$, let $g_{-1,k-1}(x)=q_{k-1}(x)$. Then
@@
$g_{k-2,k-1}(x)$ $=q_{k-1}(x)+\left(\left[\begin{array}[]{c|cccc}0&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \end{array}\right]/V_{k-2}\right)$
@@
and $\tilde{p}_{k-1}(x)=(q_{0}(x),q_{1}(x),\dots,q_{k-2}(x))V_{k-2}^{-1}(q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}}),q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}}),\dots,q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}}))^{\textup{T}}$ is precisely the interpolation polynomial which satisfies the conditions:
@@
Thus we obtain that $g_{k-2,k-1}^{(\alpha_{i})}(x_{\beta_{i}})=q_{k-1}^{(\alpha_{i})}(x_{\beta_{i}})-\tilde{p}_{k-1}^{(\alpha_{i})}(x_{\beta_{i}})=0,i=1,2,\dots,k-1.$
Next we prove the case $i=k$. The Schur complement of $V_{k-2}$ in $V_{k-1}$ is
$\left(V_{k-1}/V_{k-2}\right)$ $=\left(\left[\begin{array}[]{cccc|c}q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&&\vdots&\vdots\\ q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \hline\cr q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/V_{k-2}\right)$ (20)
$=\left(\left[\begin{array}[]{cccc|c}q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})&q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\end{array}\right]/V_{k-2}\right)$ (26)
$=\delta_{x_{\beta_{k}}}\circ D_{x}^{\alpha_{k}}\left(\left(\begin{array}[]{c|cccc}q_{k-1}(x)&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\end{array}\right)/V_{k-2}\right)$ (32)
$=\delta_{x_{\beta_{k}}}\circ D_{x}^{\alpha_{k}}(g_{k-2,k-1}(x))$
$=g_{k-2,k-1}^{(\alpha_{k})}(x_{\beta_{k}}).$
The second equality follows directly from Proposition 3, and the third equality holds due to Proposition 4. From Proposition 2 we know that
@@
Let $\textup{span}\{q_{0}(x),q_{1}(x),\dots,q_{N-1}(x)\}$ be a strongly proper interpolation space for the Birkhoff interpolation problem (2). $g_{j,k-1}(x)$, $k=1,2,\dots,N,j=0,1,\dots,k-2$ are defined by (14). Let $p_{-1}(x)=0$, then the interpolation polynomial
$p_{k-1}(x)=p_{k-2}(x)+\frac{y_{\beta_{k},\alpha_{k}}-p_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})}{g_{k-2,k-1}^{(\alpha_{k})}(x_{\beta_{k}})}g_{k-2,k-1}(x),~~k=1,2,\dots,N.$ (33)

### v2

@@ Abstract
As a generalization of Hermite interpolation problem, Birkhoff interpolation is an important subject in numerical approximation. This paper generalizes the existing Generalized Recursive Polynomial Interpolation Algorithm (GRPIA) that is used to compute the Hermite interpolation polynomial. Based on the theory of the Schur complement and the Sylvester identity, the proposed recursive algorithms are applicable to a broader class of Birkhoff interpolation problems, where each interpolation condition is given by the composition of an evaluation functional and a differential polynomial. The approach incorporates a judgment condition to ensure the problem’s well-posedness and computes a lower-degree Newton-type interpolation basis (which is also a strongly proper interpolation basis) along with the corresponding interpolation polynomial. Following the numerical examples, we analyze and compare the computational process and complexity of the proposed algorithm against traditional interpolation methods based on Gaussian elimination, and thus demonstrate that the proposed recursive approach reduces both computational cost and storage space requirements.
@@
journal: Journal of Computational and Applied Mathematics
[…]
organization=School of Mathematics and Statistics, Changchun University of Science and Technology, city=Changchun, postcode=130000, country=China
@@ 1 Introduction
Schoenberg proposed the theory of the so-called incidence matrix and its regularity in [1], after which research on the Polya’s conditions has advanced rapidly. Polya’s conditions provide a necessary and sufficient criterion for the well-posedness of a two-point interpolation problem. In [2], Turán investigated the (0,2)-interpolation problem and proposed several open problems related to Birkhoff interpolation. Shi presented a consolidation of the established results concerning the regularity of both the (0,2)-interpolation problem and specific Birkhoff interpolation problems [3]. In [4], Xia et al. investigated the linearization of univariate Birkhoff rational interpolation and demonstrated the effectiveness of these rational methods through numerical experiments. Furthermore, many researchers have investigated specific aspects of multivariate Birkhoff interpolation problems, see [5, 6, 7, 8, 9, 10].
@@
Recent work by Rhouni et al. has addressed the Birkhoff interpolation problem where the conditions involve only derivatives at the nodes. In [16], by constructing a set of high-degree Newton-type basis that can be factorized, the authors proposed the RHBPIA algorithm to determine the Birkhoff interpolation polynomial. The algorithms proposed in this paper differs from the RHBPIA in the following two aspects: 1. The proposed algorithms incorporate a core judgment condition to ensure the non-singularity of Vandermonde matrix, and then compute a lower-degree interpolation basis and the corresponding interpolation polynomial. 2. Algorithm 2 considers a more general Birkhoff interpolation problem in which each interpolation condition is given by the composition of an evaluation functional and a differential polynomial.
The structure of this paper is as follows. Section 2 introduces fundamental concepts of the Birkhoff interpolation problem, the concept of Schur complement, and the matrix Sylvester identity. Section 3 defines the concept of a strongly proper interpolation basis and derives a recursive interpolation formula for a given Birkhoff interpolation problem. Section 4 proposes two algorithms for the univariate Birkhoff interpolation problem and provides numerical examples to demonstrate their performance. The final section provides conclusions.
@@ 2 Preliminary
We first introduce some relevant concepts related to Birkhoff interpolation problem. Throughout the paper, let $\mathbb{F}$ denote either the real field $\mathbb{R}$ or the complex field $\mathbb{C}$. $\mathbb{F}[x]$ stands for the polynomial ring over $\mathbb{F}$. Let $x_{0},x_{1},\dots,x_{n}$ be $n+1$ pairwise distinct interpolation nodes. The given $N$ interpolation conditions can be expressed by linear functionals, i.e., the composite operation of evaluation functionals and differential operators
[…]
where $D_{x}=\frac{d}{dx},$ and $\alpha_{i,j}\in\mathbb{N},i=0,1,\dots,n,j=1,2,\dots,s_{i}.$ In this paper, we define the set of pairs
$\tilde{e}=\{(\beta,\alpha)\mid\beta\in\{0,1,\dots,n\},\alpha\in\{\alpha_{\beta,1},\alpha_{\beta,2},\dots,\alpha_{\beta,s_{\beta}}\}\},$ (1)
and the total number of elements in $\tilde{e}$ is $N$. The Birkhoff interpolation problem is to find a polynomial subspace $\mathcal{F}\subset\mathbb{F}[x]$ such that there exists a unique $p(x)\in\mathcal{F}$ satisfying the following interpolation conditions
@@
Here $\beta_{1},\beta_{2},\dots,\beta_{N}$ are the node numbers $0,1,\dots,n$, and $\alpha_{i},i=1,2,\dots,N$ denotes the derivative order corresponding to the interpolation node $x_{\beta_{i}}$.
In the following text, for a given Birkhoff interpolation problem, we always sort the interpolation conditions according to the above definition. The interpolation conditions are as follows:
@@
in which $T=[(\beta_{1},\alpha_{1}),(\beta_{2},\alpha_{2}),\dots,(\beta_{N},\alpha_{N})]$ is an NN-DOS. For example, let the Birkhoff interpolation problem be defined by the following conditions
$\delta_{x_{0}},\delta_{x_{1}}\circ D_{x},\delta_{x_{1}}\circ D_{x}^{2},\delta_{x_{2}}\circ D_{x}^{2},$
then, $T=[(\beta_{1},\alpha_{1}),(\beta_{2},\alpha_{2}),(\beta_{3},\alpha_{3}),(\beta_{4},\alpha_{4})]$ and $(\beta_{1},\alpha_{1})=(0,0),(\beta_{2},\alpha_{2})=(1,1),(\beta_{3},\alpha_{3})=(1,2),(\beta_{4},\alpha_{4})=(2,2)$.
@@
Next, for constructing the recursive algorithms for computing interpolation polynomials, we need to introduce the concept of Schur complement and the matrix Sylvester identity, for more details, see [19, 20, 21, 22].
@@
where the submatrix $D$ is assumed to be square and non-singular. The Schur complement of $D$ in $M$, denoted by $(M/D)$, is defined as
$(M/D)=A-BD^{-1}C.$
@@
[19] Let $D$ be a non-singular matrix and $M$ be a square matrix. Then
$|M|=|D|\cdot|(M/D)|,$
where $|\cdot|$ denotes determinant.
@@
[12, 23] Assuming that $D$ is a non-singular matrix. Then
$\left(\left(\begin{array}[]{cc}A&B\\ C&D\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}D&C\\ B&A\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}B&A\\ D&C\\ \end{array}\right)/D\right)=\left(\left(\begin{array}[]{cc}C&D\\ A&B\\ \end{array}\right)/D\right).$
@@
[12, 23] Assuming that $E$ is a given linear operator, $D$ is a non-singular matrix, and both $E\ast A$ and $E\ast B$ are well-defined, then we have
$\left(\left(\begin{array}[]{cc}E*A&E*B\\ C&D\\ \end{array}\right)/D\right)=E*\left(\left(\begin{array}[]{cc}A&B\\ C&D\\ \end{array}\right)/D\right).$
@@
([12, 22] The Sylvester identity) Let $M$ be the matrix defined in Definition 3, and $K$ be a partitioned matrix as shown below
@@
$(K/M)$ $=((K/A)/(M/A))$
$=\left(\left(\begin{array}[]{cc}E&F\\ H&A\\ \end{array}\right)/A\right)-\left(\left(\begin{array}[]{cc}F&G\\ A&B\\ \end{array}\right)/A\right)(M/A)^{-1}\left(\left(\begin{array}[]{cc}H&A\\ L&C\\ \end{array}\right)/A\right).$
@@ 3 Main results
$\mathcal{F}=\text{span}\{q_{0}(x),q_{1}(x),\dots,q_{N-1}(x)\},$
in which $\deg(q_{0}(x))<\deg(q_{1}(x))<\dots<\deg(q_{N-1}(x))$, then the interpolation polynomial $p(x)\in\mathbb{F}[x]$ can be expressed as
$p(x)=\sum_{i=0}^{N-1}a_{i}q_{i}(x),$
@@
with $Y_{N}=(y_{\beta_{1},\alpha_{1}},y_{\beta_{2},\alpha_{2}},\dots,y_{\beta_{N},\alpha_{N}})^{\textup{T}}$ and
@@
$p(x)=(q_{0}(x),q_{1}(x),\dots,q_{N-1}(x))\boldsymbol{a}=(q_{0}(x),q_{1}(x),\dots,q_{N-1}(x))V_{N-1}^{-1}Y_{N}.$
@@
where $\{q_{0}(x),q_{1}(x),\dots,q_{N-1}(x)\}$ is a strongly proper interpolation basis. This expresses $p_{N-1}(x)$ as the Schur complement of a matrix, where in Definition 3, $A=0$ is a scalar.
@@
$V_{k-1}:=\left(\begin{array}[]{ccccc}q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&&\vdots\\ q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right).$
@@
Inspired by [11, 12], we present analogous results for our problem and include proofs following its idea for completeness.
@@
[…]
@@
$g_{k-2,k-1}(x)$ $=q_{k-1}(x)$
$~~+\left(\left[\begin{array}[]{c|cccc}0&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \end{array}\right]/V_{k-2}\right)$
@@
and $\tilde{p}_{k-1}(x)=(q_{0}(x),q_{1}(x),\dots,q_{k-2}(x))V_{k-2}^{-1}(q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}}),q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}}),\dots,q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}}))^{\textup{T}}$ is precisely the interpolation polynomial which satisfies the conditions
@@
Thus we obtain $g_{k-2,k-1}^{(\alpha_{i})}(x_{\beta_{i}})=q_{k-1}^{(\alpha_{i})}(x_{\beta_{i}})-\tilde{p}_{k-1}^{(\alpha_{i})}(x_{\beta_{i}})=0,i=1,2,\dots,k-1.$
Next, we prove the case $i=k$. The Schur complement of $V_{k-2}$ in $V_{k-1}$ is
$\begin{aligned} \left(V_{k-1}/V_{k-2}\right)&=\left(\left[\begin{array}[]{cccc|c}q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&&\vdots&\vdots\\ q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \hline\cr q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/V_{k-2}\right)\\ &=\left(\left[\begin{array}[]{c|cccc}q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})&q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\end{array}\right]/V_{k-2}\right)\\ &=\delta_{x_{\beta_{k}}}\circ D_{x}^{\alpha_{k}}\left(\left(\begin{array}[]{c|cccc}q_{k-1}(x)&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots\\ q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\end{array}\right)/V_{k-2}\right)\\ &=\delta_{x_{\beta_{k}}}\circ D_{x}^{\alpha_{k}}(g_{k-2,k-1}(x))\\ &=g_{k-2,k-1}^{(\alpha_{k})}(x_{\beta_{k}}).\end{aligned}$
The second equality follows directly from Proposition 3, and the third equality is due to Proposition 4. From Proposition 2 we know that
@@
[…]
$p_{k-1}(x)=p_{k-2}(x)+\frac{y_{\beta_{k},\alpha_{k}}-p_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})}{g_{k-2,k-1}^{(\alpha_{k})}(x_{\beta_{k}})}g_{k-2,k-1}(x),~~k=1,2,\dots,N.$